### v2

# Weak type (1,1) jump inequalities in a nonsymmetric Gaussian setting

Abstract.
We prove that the jump quasi-seminorm of order $\varrho=2$ for a general Ornstein–Uhlenbeck semigroup $\left(\mathcal{H}_{t}\right)_{t>0}$ in $\mathbb{R}^{n}$ defines an operator of weak type $(1,1)$ with respect to the invariant measure. This provides an example of a weak-type jump inequality for a nonsymmetric semigroup in a nondoubling measure space.
Our result may be seen as an endpoint refinement of the weak type $(1,1)$ inequality for the $\varrho$-th order variation seminorm of $\left(\mathcal{H}_{t}\right)_{t>0}$, recently proved by the authors when $\varrho>2$, and disproved for $\varrho=2$.
Key words and phrases:
Jump quasi-seminorm, variation seminorm, jump inequalities, Ornstein–Uhlenbeck semigroup,Mehler kernel
2020 Mathematics Subject Classification:
37A46, 37A30,47D03, 42B99
The first and second authors are members of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). This research was carried out while the third author was visiting the University of Padova, Italy, and he is grateful for its hospitality. The third author also profited from a grant from the Royal Swedish Academy of Sciences and another from the Royal Society of Arts and Sciences in Gothenburg.
1. Introduction
In this paper we study jump and variational bounds of weak type $(1,1)$ for a nonsymmetric Ornstein–Uhlenbeck semigroup, that is, for the semigroup $(\mathcal{H}_{t})_{t\geq 0}$ generated by the operator
$\mathcal{L}f(x)=\frac{1}{2}\,\operatorname{tr}\big(Q\nabla^{2}f\big)(x)+\langle Bx,\nabla f(x)\rangle\,,\qquad f\in\mathcal{S}(\mathbb{R}^{n}),\;\;x\in\mathbb{R}^{n}.$
Here $\nabla$ is the gradient and $\nabla^{2}$ the Hessian, and $n\geq 1$. The definition of $\mathcal{L}$ depends on two real $n\times n$ matrices, $Q$ and $B$; we assume that $Q$ is symmetric and positive definite, and that all the eigenvalues of $B$ have negative real parts. In the following, $\mathbb{R}^{n}$ is equipped with a Gaussian measure denoted by $\gamma_{\infty}$, which is invariant under the action of $\mathcal{H}_{t}$; see (3.7) for a precise definition of $\gamma_{\infty}$ in terms of $Q$ and $B$. This measure is locally but not globally doubling. Under our assumptions on $Q$ and $B$, neither $\mathcal{L}$ nor $\mathcal{H}_{t}$ is symmetric in $L^{2}(\gamma_{\infty})$ in general.
[…]
$\|\mathcal{H}_{t}\,f(x)\|_{v(\varrho),\mathcal{I}}=\sup\,\left(\sum_{i=1}^{N}|\mathcal{H}_{t_{i}}f(x)-\mathcal{H}_{t_{i-1}}f(x)|^{\varrho}\right)^{1/\varrho},\qquad x\in\mathbb{R}^{n},\quad 1\leq\varrho<\infty,$ (1.1)
the supremum being over all finite, increasing sequences $\left(t_{i}\right)_{0}^{N}$ of points in $\mathcal{I}$. We emphasize that in this paper the $v(\varrho)$ seminorm will be always taken in the variable $t$.
In a recent paper the authors proved that for $\varrho>2$ the variation operator mapping $f\in L^{1}(\gamma_{\infty})$ to the function
$x\mapsto\|\mathcal{H}_{t}f(x)\|_{v(\varrho),\mathbb{R}_{+}}$
is bounded from $L^{1}(\gamma_{\infty})$ to $L^{1,\infty}(\gamma_{\infty})$ (see [15], and [14] for a different proof in the one-dimensional case). The boundedness of this variation operator on $L^{p}(\gamma_{\infty}),\;1<p<\infty,$ was already known (see [26, Corollary 4.5], and also [2]).
These results break down for $1\leq\varrho\leq 2$, as proved in [15, Section 8].
Our main result in this paper is a jump inequality of weak type $(1,1)$ for $\mathcal{H}_{t}$ at $\varrho=2$. We therefore need to introduce jump quasi-seminorms, which we do only in the Ornstein–Uhlenbeck framework.
[…]
Then the jump quasi-seminorm of order $\varrho$ of $\mathcal{H}_{t}\,f$ in $\mathcal{I}$ is
$J_{\varrho}^{p}(\mathcal{H}_{t}f:\,t\in\mathcal{I})=\sup_{\lambda>0}\|\lambda\,\big(N_{\lambda}\big(\mathcal{H}_{t}f(\cdot):t\in\mathcal{I}\big)\big)^{1/\varrho}\|_{L^{p}(\gamma_{\infty})},\quad 1\leq p,\varrho<\infty.$ (1.3)
Notice that the quasi-seminorms $J_{\varrho}^{p}(\cdot)$ are decreasing in $\varrho$.
One can define weak jump quasi-seminorms in a similar way, by replacing the Lebesgue space $L^{p}(\gamma_{\infty})$ in (1.3) by a Lorentz space $L^{p,q}(\gamma_{\infty})$ (see [31, formula (1.2)]); we are mainly interested in the case $(p,q)=(1,\infty)$, where
$J_{\varrho}^{1,\infty}(\mathcal{H}_{t}f:\,t\in\mathcal{I})=\sup_{\lambda>0}\|\lambda\,\big(N_{\lambda}\big(\mathcal{H}_{t}f(\cdot):t\in\mathcal{I}\big)\big)^{1/\varrho}\|_{L^{1,\infty}(\gamma_{\infty})}.$ (1.4)
Since
$\lambda\big(N_{\lambda}(\mathcal{H}_{t}f(x):\,\,t\in\mathcal{I})\big)^{1/\varrho}\leq\|\mathcal{H}_{t}\,f(x)\|_{v(\varrho),\mathcal{I}}\,,\qquad x\in\mathbb{R}^{n},$
for all $\varrho\geq 1$ and $\lambda>0$ (see [31, formula (2.3)]), the jump quasi-seminorms $J_{\varrho}^{p}$ are dominated by the $L^{p}$ norms of variational seminorms of order $\varrho$, that is,
$J_{\varrho}^{p}(\mathcal{H}_{t}f:\,t\in\mathcal{I})\leq\|\,\|\mathcal{H}_{t}\,f(\cdot)\|_{v(\varrho),\mathcal{I}}\|_{L^{p}(\gamma_{\infty})},\qquad\varrho\geq 1,\quad p\geq 1.$ (1.5)
It also follows that
$\big|\{x\in\mathbb{R}^{n}$ $:\lambda\left(N_{\lambda}(\mathcal{H}_{t}f(x):\,t\in\mathcal{I})\right)^{1/\varrho}>\alpha\}\big|$
$\leq\left|\{x\in\mathbb{R}^{n}:\|\mathcal{H}_{t}f(x)\|_{v(\varrho),\mathcal{I}}>\alpha\}\right|,\qquad\alpha>0,$
and thus
$J_{\varrho}^{1,\infty}(\mathcal{H}_{t}f:\,t\in\mathcal{I})\leq\|\,\|\mathcal{H}_{t}f(\cdot)\|_{v(\varrho),\mathcal{I}}\|_{L^{1,\infty}(\gamma_{\infty})},\qquad\varrho\geq 1.$ (1.6)
Notice that combining (1.5) with the results in [26, Corollary 4.5] and [2, p. 31] on the strong boundedness $L^{p}-L^{p}$ of the operator $f\mapsto\|\mathcal{H}_{t}f\|_{v(\varrho),\mathbb{R}_{+}}$, for $1<p<\infty$ and $\varrho>2$ one immediately arrives at the following result.
Proposition 1.1.
Let $1<p<\infty$ and $\varrho>2$. Then the jump quasi-seminorms of the Ornstein–Uhlenbeck semigroup $\left(\mathcal{H}_{t}\right)_{t>0}$ are bounded on $L^{p}(\gamma_{\infty})$, that is,
[…]
On the other hand, for $\varrho>2$ one has
$\|\,\|\mathcal{H}_{t}\,f\|_{v(\varrho),\mathcal{I}}\|_{L^{1,\infty}(\gamma_{\infty})}\lesssim_{\varrho}J_{2}^{1}(\mathcal{H}_{t}\,f:\,t\in\mathcal{I})\leq\|\,\|\mathcal{H}_{t}\,f\|_{v(2),\mathcal{I}}\|_{L^{1}(\gamma_{\infty})},$ (1.7)
with the first inequality failing for $\varrho=2$. See [32, formula (1.4)] or [31, Lemma 2.3] for the first inequality; the second is (1.5) with $p=1$. Thus a jump inequality of weak type $(1,1)$ for $\mathcal{H}_{t}\,f$ at $\varrho=2$ can be considered an endpoint refinement of the variational estimate.
We recall that jump and variational inequalities say that the jump quasi-seminorms or the $L^{p}$ norms of the variation seminorms of $(\mathcal{H}_{t}f)_{t\in\mathcal{I}}$ are bounded by the $L^{p}$ norm of $f$. Since the pioneering works of Lépingle, Gaposhkin and Bourgain [27, 17, 18, 7, 8], this kind of bounds has proved to be very useful to measure and control the fluctuations of families of linear operators, and it often provides an answer to the problem of pointwise convergence. The range of operators for which variational bounds have been proved in the last thirty years is vaste (see e.g. [20, 9, 21, 22, 24, 19, 23, 16, 4, 29, 30, 28, 31, 32, 33, 3, 5, 34, 35, 36, 6]). But only a few results seem to concern the case of an Ornstein–Uhlenbeck semigroup, despite the prototypical nature of $(\mathcal{H}_{t})_{t\geq 0}$ as a possibly nonsymmetric semigroup in a nondoubling measure space.
The standard case of a symmetric Ornstein-Uhlenbeck semigroup is given by $Q=-B=I_{n}$, where $I_{n}$ is the $n$-dimensional identity matrix. In that particular case, a variational approach was successfully addressed in [19, 16], for $\varrho>2$. Recent contributions in the nonsymmetric Ornstein–Uhlenbeck context are due to Almeida, Betancor, Farina, Quijano and Rodriguez-Mesa [1, 2]; they prove some interesting variational inequalities, under the condition $\varrho>2$.
The counterexample exhibited in [15, Section 8], which is based on earlier work due to Qian [39], says that for $\varrho=2$ the variation seminorm of $\left(\mathcal{H}_{t}\right)_{t>0}$ is neither of strong nor of weak type $(p,p)$ for any $p\in[1,\infty)$, with respect to $\gamma_{\infty}$. In the light of this counterexample and of (1.5) and (1.6), it is reasonable to look for a weak type $(1,1)$ jump inequality at $\varrho=2$. Our main result is the following.
Theorem 1.2.
For all $f\in L^{1}(\gamma_{\infty})$ one has
[…]
with $C=C(n,Q,B)$.
For $\varrho<2$ an analogous estimate is not possible. Indeed, it has been proved in [31, Lemma 2.3] and in [34, formula (2.20)], both applied with $r=2$ and $p=1$, that for all $1\leq\varrho<2$ one has
$\|\,\|\mathcal{H}_{t}\,f(x)\|_{v(2),\mathcal{I}}\|_{L^{1,\infty}(\gamma_{\infty})}\lesssim_{\varrho}J_{\varrho}^{1,\infty}(\mathcal{H}_{t}\,f:\,t\in\mathcal{I}).$ (1.9)
As a consequence, the weak jump quasi-seminorms $J_{\varrho}^{1,\infty}(\mathcal{H}_{t}\,f:\,t\in\mathcal{I})$ are unbounded for all $\varrho\in[1,2)$. We remark that in (1.9) one cannot replace $L^{1,\infty}(\gamma_{\infty})$ by $L^{1}(\gamma_{\infty})$, as proved in [35, Lemma 2.5, p. 3395]; this lemma is stated with $\varrho=2$, but the proof holds for any $1\leq\varrho\leq 2$.
In order better to understand the scheme of the proof of Theorem 1.2 in the following sections, we need to go a little deeper into the proof of [15, Theorem 1.1], concerning the weak type $(1,1)$ of the $\varrho$-th order variation seminorm of $\left(\mathcal{H}_{t}\right)_{t>0}$. A basic ingredient of our analysis was the Mehler kernel $K_{t}=K_{t}(x,u)$, that is, the integral kernel of the semigroup operator $\mathcal{H}_{t}$ with respect to the invariant measure; see (4.2) for an explicit expression of $K_{t}$. It is well-known that $K_{t}$ behaves very differently in the global and in the local region, that is, roughly speaking, away from or close to the diagonal $x=u$ (more precise definitions will be given in Section 5). It also matters whether the time parameter $t$ is large or small. For this reason, the proof of [15, Theorem 1.1] requires several distinctions both in space and in time.
Some parts of the operator $f\mapsto\|\mathcal{H}_{t}f\|_{v(\varrho),\mathbb{R}_{+}}$ turn out to be of weak type $(1,1)$ not only for $\varrho>2$, but even for $\varrho\geq 1$. Because of (1.6), the corresponding parts of $J_{\varrho}^{1,\infty}(\mathcal{H}_{t}f:\,t>0)$ will then be controlled by the $L^{1}(\gamma_{\infty})$ norm of $f$. In this way, we shall see in Section 5 how the proof of Theorem 1.2 can be reduced to the case of the local region for $t\in(0,1]$. This case will be treated in the last four sections, where we compare our semigroup with a convolution semigroup (essentially the heat semigroup), for which jump bounds are known [28]. It is worth mentioning that, differently from [15], we prove Theorem 1.2 without appealing to vector-valued Calderón–Zygmund theory.
Unexpectedly, the question whether the jump quasi-seminorm of a general Ornstein–Uhlenbeck semigroup in $\mathbb{R}^{n}$ defines an operator of strong type $(p,p),\;1<p<\infty,$ with respect to the invariant measure, is still open. Notice however that in the symmetric case (that is, when each operator $\mathcal{H}_{t}$ is a self-adjoint operator) the boundedness of $J^{p}_{2}(\mathcal{H}_{t}f:t>0)$ follows from [31, Theorem 1.2].
The structure of the article is as follows. Section 2 gives a subadditivity property for the counting function. Then in Sections 3 and 4 we state basic facts concerning the Ornstein–Uhlenbeck semigroup and the Mehler kernel, respectively. The distinction between the local and the global part of the Ornstein–Uhlenbeck semigroup operator is presented in Section 5, and some cases of Theorem 1.2 are verified (see above). The remaining case of Theorem 1.2 is proved in Sections 6–9. This is the local part for $t\in(0,1]$, and the localization allows us to deal with one patch of $\mathbb{R}^{n}$ at a time. More precisely, in Section 6 we treat the case when $t\leq 1$ is not too small (essentially $|x|^{-2}\leq t\leq 1$), proving the strong type $(p,p)$ for $p>1$ and the weak type $(1,1)$ for the variation of the local part of $\mathcal{H}_{t}f$, and here $\varrho\geq 1$. Dealing with the local part of $\mathcal{H}_{t}f$ for small $t$ (essentially $0<t\leq\min(1,{|x|^{-2}})$) is more involved. In Section 7 we decompose the semigroup operator into a sum of three difference operators and a main operator. By means of some technical estimates one proves in Section 8 strong and weak type estimates for the variation of the difference operators, with $\varrho\geq 1$. In Section 9 we conclude the proof of Theorem 1.2 by showing that the main operator satisfies both strong and weak jump inequalities for $\varrho=2$.
It is a pleasure to thank Professor Mariusz Mirek for many helpful and constructive discussions on the subject of this paper.
Notation
We will write $C<\infty$ and $c>0$ for various constants, all of which depend only on $n$, $Q$ and $B$, unless otherwise explicitly stated. The constants involved when we write $\lesssim$ or $\gtrsim$ or $\simeq$ are of this type, except when there is a subscript indicating dependence also on an additional parameter. The Lebesgue measure of a measurable set $E\subset\mathbb{R}^{n}$ will be denoted by $|E|$. We write $B(x_{0},r)$ for the closed ball of center $x_{0}$ and radius $r$, and $CB(x_{0},r)$ for its concentric scaling. By $\mathbb{N}$ we mean $\{0,1,\dots\}$.
2. Subadditivity
In this section, we describe a subadditivity property of the counting function, in a generic setting. For a more comprehensive treatment of the variational and jump inequalities, we refer the reader to [34, 35] or the recent monograph [25].
[…]
$N_{\lambda}\big(\varphi(t)+\psi(t):\,t\in\mathcal{I}\big)\leq 2\Big(N_{\lambda/4}\big(\varphi(t):\,t\in\mathcal{I}\big)+N_{\lambda/4}\big(\psi(t):\,t\in\mathcal{I}\big)\Big)$ (2.1)
for all $\varphi,\psi$ defined on $\mathcal{I}$ and all $\lambda>0$. For a proof we refer to [23, formulas (2) and (3), p. 6712], where a slightly different $\lambda$-jump counting function is also involved. This justifies calling $J_{\varrho}^{p}(\cdot)$ and $J_{\varrho}^{1,\infty}(\cdot)$ quasi-seminorms.
3. A general Gaussian setting
We introduce some basic notions in the setting of Ornstein–Uhlenbeck analysis and recall some facts from [10, 11, 12].
  1. (1)
    As stated in the introduction, $Q$ is a real, symmetric and positive definite $n\times n$ matrix, and $B$ a real $n\times n$ matrix whose eigenvalues have negative real parts.
  2. (2)
    The covariance matrices are defined by
    $Q_{t}=\int_{0}^{t}e^{sB}Qe^{sB^{*}}ds,\qquad\text{ $t\in(0,+\infty]$}.$ (3.1)
    Both $Q_{t}$ and $Q_{\infty}$ are well defined, symmetric and positive definite. Since $Q_{t}=t\,Q(1+\mathcal{O}(t))$ as $t\to 0$, one also has
    $Q_{t}^{-1}=t^{-1}\,Q^{-1}(1+\mathcal{O}(t))$ (3.2)
    and
    $\det\,Q_{t}=t^{n}\,\det Q\,(1+\mathcal{O}(t))$ (3.3)
    as $t\to 0$. Further,
    $|Q_{t}^{-1}|\lesssim t^{-1}\quad\mathrm{for}\quad 0<t<1.$ (3.4)
  3. (3)
    $R(x)$ denotes the quadratic form
    $R(x)={\frac{1}{2}\left\langle Q_{\infty}^{-1}x,x\right\rangle},\qquad\text{$x\in\mathbb{R}^{n}$.}$
  4. (4)
    $D_{t}$ is a one-parameter semigroup of matrices, defined by
    $D_{t}=Q_{\infty}e^{-tB^{*}}Q_{\infty}^{-1},\qquad t\in\mathbb{R}.$ (3.5)
    For $|t|\leq 1$ and $0\neq x\in\mathbb{R}^{n}$ these matrices satisfy
    $|x-D_{t}x|\simeq|t|\,|x|,$ (3.6)
  5. (5)
    $\gamma_{t}$ denotes the normalized Gaussian measure in $\mathbb{R}^{n}$, given by
    $d\gamma_{t}(x)=(2\pi)^{-\frac{n}{2}}(\text{det}\,Q_{t})^{-\frac{1}{2}}e^{-\frac{1}{2}\langle Q_{t}^{-1}x,x\rangle}dx\,,\qquad\text{ $t\in(0,+\infty]$. }$
    In particular,
    $d\gamma_{\infty}(x)=(2\pi)^{-\frac{n}{2}}(\text{det}\,Q_{\infty})^{-\frac{1}{2}}e^{-R(x)}dx,$ (3.7)
    and this is the unique invariant measure for the Ornstein–Uhlenbeck semigroup.
4. The Mehler kernel and its derivative
The Ornstein–Uhlenbeck semigroup operator $\mathcal{H}_{t}$ is given, for all $f\in L^{1}(\gamma_{\infty})$ and all $t>0$, by
$\mathcal{H}_{t}f(x)$ $=\int K_{t}(x,u)\,f(u)\,d\gamma_{\infty}(u)\,,\qquad x\in\mathbb{R}^{n}.$ (4.1)
The integral kernel $K_{t}$ is called the Mehler kernel and is given, for all $x,u\in\mathbb{R}^{n}$ and $t>0$, by
$K_{t}(x,u)\!=\!\Big(\frac{\det\,Q_{\infty}}{\det\,Q_{t}}\Big)^{{1}/{2}}e^{R(x)}\exp\Big[{-\frac{1}{2}\left\langle\left(Q_{t}^{-1}-Q_{\infty}^{-1}\right)(u-D_{t}\,x),u-D_{t}\,x\right\rangle}\Big],\qquad$ (4.2)
see [10, formula (2.6)].
The following estimate, holding for $0<t\leq 1$ and all $x,u\in\mathbb{R}^{n}$, is proved in [10, (3.4)]:
$\frac{e^{R(x)}}{t^{n/2}}\exp\left[-C\,\frac{|u-D_{t}\,x|^{2}}{t}\right]\,\lesssim\,K_{t}(x,u)\,\lesssim\,\frac{e^{R(x)}}{t^{n/2}}\exp\left[-c\,\frac{|u-D_{t}\,x|^{2}}{t}\right].$ (4.3)
We shall also need the derivative of $K_{t}(x,u)$ with respect to $t$, which we denote by $\dot{K}_{t}(x,u)$. For all $(x,u)\in\mathbb{R}^{n}\times\mathbb{R}^{n}$ and $t>0$, we proved in [13, Lemma 4.2] that
$\dot{K}_{t}(x,u)=K_{t}(x,u)\,N_{t}^{(0)}(x,u),$ (4.4)
where the function $N_{t}^{(0)}$ is given by
$N_{t}^{(0)}(x,u)$ $=-\frac{1}{2}\,{\operatorname{tr}\big(Q_{t}^{-1}\,e^{tB}\,Q\,e^{tB^{*}}\big)}+\frac{1}{2}\,\left|Q^{1/2}\,e^{tB^{*}}\,Q_{t}^{-1}\,(u-D_{t}\,x)\right|^{2}$
$\qquad\qquad\qquad-\left\langle Q_{\infty}\,B^{*}\,Q_{\infty}^{-1}\,D_{t}\,x\,,\,(Q_{t}^{-1}-Q_{\infty}^{-1})\,(u-D_{t}\,x)\right\rangle.$ (4.5)
For future convenience, we describe in the Ornstein–Uhlenbeck framework a simple way of estimating the variation seminorm, which is sometimes applicable. In the case when $\varrho=1$, the variation of $\mathcal{H}_{t}f(x)$ is a supremum of sums $\sum|\mathcal{H}_{t_{i}}f(x)-\mathcal{H}_{t_{i-1}}f(x)|$, which can be estimated by the sum $\sum\int_{t_{i-1}}^{t_{i}}|\partial_{t}\mathcal{H}_{t}f(x)|\,dt$. Since the variation seminorm is decreasing in $\varrho$, we get for each $\varrho\in[1,\infty)$ and any interval $I\subset\mathbb{R}_{+}$, at least formally,
$\|\mathcal{H}_{t}f(x)\|_{v(\varrho),I}$ $\leq\|\mathcal{H}_{t}f(x)\|_{v(1),I}\leq\int_{I}\left|\frac{\partial}{\partial t}\int K_{t}(x,u)f(u)\,d\gamma_{\infty}(u)\right|\,dt$ (4.6)
$\leq\int\int_{I}\big|\dot{K}_{t}(x,u)\big|\,dt\,|f(u)|\,d\gamma_{\infty}(u),\qquad f\in L^{1}(\gamma_{\infty}).$
This argument was justified and used in [13, Subsection 5.1]. Below we will apply it to some other operators and their kernels.
5. Localization
The distinction between a local and a global part of the operator is a standard tool in the study of the Ornstein–Uhlenbeck semigroup; we refer to [37, 40, 38] for more details on this idea. Here we just describe in a concise way the localization procedure introduced in [13, Section 7].
We choose a sequence $B_{j}=B(x_{j},1/(1+|x_{j}|)),\;j=0,1,\dots,$ of closed balls whose interiors are pairwise disjoint, and they form a maximal family with this property. As shown in [13, Section 7], the scaled balls $3B_{j}=B(x_{j},3/(1+|x_{j}|))$ then cover $\mathbb{R}^{n}$. The sequence starts with $B_{0}=B(0,1)$, which implies that $|x_{j}|>1$ for $j\geq 1$.
In [13, Section 7] we also introduced smooth, nonnegative funtions $r_{j},\;j=0,1,\dots,$ such that $\sum_{0}^{\infty}r_{j}=1$ in $\mathbb{R}^{n}$, and with $r_{j}$ supported in $4B_{j}$. The functions $\widetilde{r}_{j},\;j=0,1,\dots,$ also have values in $[0,1]$, and $\widetilde{r}_{j}$ is supported in $6B_{j}$ and equals 1 in $5B_{j}$. For the gradients, one has
$|\nabla r_{j}(x)|,\;\;\big|\nabla\widetilde{r}_{j}(x)\big|\lesssim 1+|x|.$ (5.1)
The balls $6B_{j}$ have bounded overlap, see [13, Section 7]. Moreover, the density of $\gamma_{\infty}$ is essentially constant in each $6B_{j}$, since
$e^{R(x)}\simeq e^{R(x_{j})},\qquad x\in 6B_{j},$ (5.2)
uniformly in $j$.
Then we let
$\eta(x,u)=\sum_{j=0}^{\infty}\widetilde{r}_{j}(x)\,r_{j}(u),\qquad x,\;u\in\mathbb{R}^{n},$ (5.3)
and observe that $0\leq\eta(x,u)\leq 1.$ The local and global parts of $\mathcal{H}_{t}$ are defined as
$\mathcal{H}_{t}^{\mathrm{loc}}f(x)=\int_{\mathbb{R}^{n}}K_{t}(x,u)\,\eta(x,u)\,f(u)\,d\gamma_{\infty}(u),\qquad x\in\mathbb{R}^{n},$ (5.4)
and
$\mathcal{H}_{t}^{\mathrm{glob}}f(x)=\mathcal{H}_{t}f(x)-\mathcal{H}_{t}^{\mathrm{loc}}f(x)=\int_{\mathbb{R}^{n}}K_{t}(x,u)\,(1-\eta(x,u))\,f(u)\,d\gamma_{\infty}(u),\qquad x\in\mathbb{R}^{n}.$
We observe that
$\mathcal{H}_{t}^{\mathrm{loc}}f(x)=\sum_{0}^{\infty}\widetilde{r}_{j}(x)\,\mathcal{H}_{t}(f\,r_{j}),\qquad x\in\mathbb{R}^{n}.$ (5.5)
This sum is locally finite, since the $\widetilde{r}_{j}$ have bounded overlap.
Some parts of our jump operator are easy to treat, by means of the following proposition.
Proposition 5.1.
Let $1\leq\varrho<\infty$. For $f\in L^{1}(\gamma_{\infty})$
$J^{1,\infty}_{\varrho}\left(\mathcal{H}_{t}f:t\in[1,\infty)\right)\lesssim_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varrho}\|f\|_{L^{1}(\gamma_{\infty})}$ (5.6)
and
$J^{1,\infty}_{\varrho}\left(\mathcal{H}_{t}^{\mathrm{glob}}f:t\in(0,1]\right)\lesssim_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varrho}\|f\|_{L^{1}(\gamma_{\infty})}.$ (5.7)
Proof.
From [15, Theorems 4.3 and 6.1] we know that for $1\leq\varrho<\infty$ the operators mapping $f\in L^{1}(\gamma_{\infty})$ to the functions
$\|\mathcal{H}_{t}f(x)\|_{v(\varrho),[1,+\infty)},\quad x\in\mathbb{R}^{n},$ (5.8)
and
$\|{\mathcal{H}}_{t}^{\mathrm{glob}}f(x)\|_{v(\varrho),(0,1]},\quad x\in\mathbb{R}^{n},$ (5.9)
are of weak type $(1,1)$ with respect to the measure $\gamma_{\infty}$. (Notice that our definition of ${\mathcal{H}}_{t}^{\mathrm{glob}}$ is that of [13] so the maximal operator estimate used in the proof of [15, Theorem 6.1] is that of [13, Theorem 10.1], not [15, Theorem 6.2].) From these two estimates and (1.6), the proposition follows. ∎
As a consequence, in order to prove Theorem 1.2 it suffices to consider the local part of the semigroup with $t\in(0,1]$. Doing so, we deal with each term in the sum in (5.5) separately, in the following way.
Proposition 5.2.
Let $j\in\mathbb{N}$. For any $f\in L^{1}(\gamma_{\infty})$
$J_{2}^{1,\infty}\left(\widetilde{r}_{j}\,\mathcal{H}_{t}(f\,r_{j}):\,0<t\leq 1\right)\lesssim\|f\|_{L^{1}(\gamma_{\infty})}$
uniformly in $j$.
Notice that (5.2) allows us to replace $\gamma_{\infty}$ here by Lebesgue measure.
Because of the bounded overlap of the $\widetilde{r}_{j}$, this proposition implies that
$J_{2}^{1,\infty}\left(\mathcal{H}_{t}^{\mathrm{loc}}f:\,0<t\leq 1\right)\lesssim\|f\|_{L^{1}(\gamma_{\infty})}.$
This and Proposition 5.7 will complete the proof of Theorem 1.2, in view of the quasi-subadditivity in Section 2.
To prove Proposition 5.2, we will split the $t$ interval $(0,1]$ for $j\geq 1$ into $|x_{j}|^{-2},1]$ (treated in Section 6) and $(0,\min(1,|x_{j}|^{-2})$ (in Sections 7, 8 and 9). For $j=0$ only the second of these subintervals will be nonempty, since $x_{0}=0$.
6. The local case for $t$ not too small
Proposition 6.1.
Let $1\leq\varrho<\infty$. For all $j\geq 1$ the operator that maps $f$ to the function
$\|\widetilde{r}_{j}\,\mathcal{H}_{t}(f\,r_{j})\|_{v(\varrho),[|x_{j}|^{-2},1]}$
is of strong type $(p,p)$ for all $p\geq 1$ with respect to Lebesgue measure, uniformly in $j$.
Because of (1.6), this implies that part of Proposition 5.2 which deals with $t\in[|x_{j}|^{-2},1]$.
Proof.
We fix $j\geq 1$. From (4.3) it follows that for $x\in 6B_{j}$, any $u\in\mathbb{R}^{n}$ and $|x_{j}|^{-2}\leq t\leq 1$
$K_{t}(x,u)$ $\lesssim e^{R(x)}\,t^{-n/2}\,\lesssim\,e^{R(x_{j})}\,|x_{j}|^{n}.$ (6.1)
Let $N(x,u)$ be the number of zeros in $|x_{j}|^{-2},1]$ of the function $t\mapsto\dot{K}_{t}(x,u)$. Proposition 9.1 in [13] implies that $N(x,u)\leq\bar{N}$ for some constant $\bar{N}$ that depends only on $n$ and $B$. We denote these zeros by $t_{1}(x,u)<\dots<t_{N(x,u)}(x,u)$, and set $t_{0}(x,u)=|x_{j}|^{-2}$ and $t_{N(x,u)+1}(x,u)=1$. It follows from the fundamental theorem of calculus that
$\int_{|x_{j}|^{-2}}^{1}\left|\dot{K}_{t}(x,u)\right|dt=\sum_{i=0}^{N(x,u)}\left|K_{t_{i+1}(x,u)}(x,u)-K_{t_{i}(x,u)}(x,u)\right|\lesssim\bar{N}\sup_{|x_{j}|^{-2}\leq t\leq 1}K_{t}(x,u).$ (6.2)
We can safely apply (4.6) to the operator $f\mapsto\widetilde{r}_{j}\,\mathcal{H}_{t}\,(f\,r_{j})$. This was verified for $\mathcal{H}_{t}$ in [13, Subsection 5.1], and the extra factors $\widetilde{r}_{j}$ and $r_{j}$ do not matter here. Combined with (6.2) and (6.1), this yields for any $\varrho\geq 1$ and $f\in L^{1}(\gamma_{\infty})$
$\|\widetilde{r}_{j}(x)\,\mathcal{H}_{t}\,(f\,r_{j})(x)\|_{v(\varrho),[|x_{j}|^{-2},1]}$ $\leq\widetilde{r}_{j}(x)\,\int\,\int_{|x_{j}|^{-2}}^{1}\big|\dot{K}(x,u)\big|\,dt\,|f(u)|\,r_{j}(u)\,d\gamma_{\infty}(u)$ (6.3)
$\leq\bar{N}\int_{4B_{j}}\,\sup_{t\in[|x_{j}|^{-2},1]}|K_{t}(x,u)|\,|f(u)|\,d\gamma_{\infty}(u)\,\mathbbm{1}_{6B_{j}}(x)$
$\lesssim\bar{N}\,e^{R(x_{j})}\,|x_{j}|^{n}\,\|f\|_{L^{1}(4B_{j},\,\gamma_{\infty})}\,\mathbbm{1}_{6B_{j}}(x)$ (6.4)
$\lesssim\big|4B_{j}\big|^{-1}\|f\|_{L^{1}(4B_{j},du)}\,\mathbbm{1}_{6B_{j}}(x).$ (6.5)
The last expression is a simple mean value of $|f|$, and the strong type $(p,p)$ for all $p\geq 1$ is clear. ∎
7. The local case for small $t$ : the decomposition
In this and the following sections, we keep $j\in\mathbb{N}$ fixed and prove Proposition 5.2 in the remaining case $0<t\leq\min(1,|x_{j}|^{-2})$. For these $t$ one has $t\lesssim 1/|x|^{2}$, as seen by considering small and large $x_{j}$ separately.
The operator $\mathcal{H}_{t}^{\mathrm{loc}}$ will be decomposed into a sum of three difference operators and a main operator. More precisely, we write
$\mathcal{H}_{t}^{\mathrm{loc}}f(x)=\sum_{\kappa=1}^{3}\Delta_{t}^{(\kappa)}f(x)+M_{t}f(x),$ (7.1)
where
$\Delta_{t}^{(\kappa)}f(x)$ $=\int_{\mathbb{R}^{n}}\Big(\widetilde{K}_{t}^{(\kappa-1)}(x,u)-\widetilde{K}_{t}^{(\kappa)}(x,u)\Big)\,f(u)\,d\gamma_{\infty}(u),\qquad\kappa=1,2,3,\qquad$ (7.2)
$M_{t}f(x)$ $=\int_{\mathbb{R}^{n}}\widetilde{K}_{t}^{(3)}(x,u)\,f(u)\,d\gamma_{\infty}(u),$ (7.3)
and the kernels $\widetilde{K}_{t}^{(\kappa)}(x,u)$,  $\kappa=1,2,3$, are given, respectively, by
$\widetilde{K}_{t}^{(0)}(x,u)$ $=K_{t}(x,u)\,$
$\widetilde{K}_{t}^{(1)}(x,u)$ $=\Big(\frac{\det Q_{\infty}}{\det Q_{t}}\Big)^{1/2}\,e^{R(x)}\,\exp\Big(-\frac{1}{2}\,\langle(Q_{t}^{-1}-Q_{\infty}^{-1})(u-x),\,u-x\rangle\Big),\qquad$ (7.4)
$\widetilde{K}_{t}^{(2)}(x,u)$ $=\Big(\frac{\det Q_{\infty}}{\det Q_{t}}\Big)^{1/2}\,e^{R(x)}\,\exp\left(-\frac{1}{2t}\,\big|Q^{-1/2}(u-x)\big|^{2}\right),$ (7.5)
$\widetilde{K}_{t}^{(3)}(x,u)$ $=\frac{(\det Q_{\infty})^{1/2}}{{(\det Q)}^{1/2}\,t^{n/2}}\,e^{R(x)}\,\exp\left(-\frac{1}{2t}\,\big|Q^{-1/2}(u-x)\big|^{2}\right).$ (7.6)
In the next section, we shall see that the three difference operators satisfy good variational inequalities for all $\varrho\geq 1$. Thus the critical part of $\mathcal{H}_{t}^{\mathrm{loc}}$ or $\mathcal{H}_{t}$ is $M_{t}$. One notices that $M_{t}$ basically coincides, up to some constants and the factor $e^{R(x)}$, with the operator with kernel $K_{t}^{c}(x-u)$ introduced in [15, formula (8.3)]. The results in [15, Section 8] therefore imply that the variation operator
$f\mapsto\|M_{t}f\|_{v(2),\mathbb{R}_{+}}$
is not of strong nor weak type $(p,p)$ with respect to $\gamma_{\infty}$, for any $p\in[1,\infty)$. It follows that the same negative result holds for $\mathcal{H}_{t}^{\mathrm{loc}}$ and $\mathcal{H}_{t}$.
However, we shall prove in Section 9 that jump quasi-seminorms with $\varrho=2$ behave well for $M_{t}$, in the way stated in Theorem 1.2.
Before considering $M_{t}$, we study the three difference operators.
8. The local case for small $t$ : the difference operators
In order to estimate the variation of the difference operators, we shall apply (4.6) to each $\Delta_{t}^{(\kappa)}$.
The necessary bounds on the $t$ derivatives of the kernels of $\Delta_{t}^{(\kappa)}$ will be proved in the following subsection.
8.1. Some auxiliary bounds
First of all, we provide expressions for $\partial_{t}\,\widetilde{K}_{t}^{(\kappa)}$, similar to (4.4).
Lemma 8.1.
For all $(x,u)\in\mathbb{R}^{n}\times\mathbb{R}^{n}$,  $t>0$ and $\kappa=1,2,3$, we have
$\partial_{t}\,\widetilde{K}_{t}^{(\kappa)}(x,u)=\widetilde{K}_{t}^{(\kappa)}(x,u)\,N_{t}^{(\kappa)}(x,u),$ (8.1)
where
$N_{t}^{(1)}(x,u)$ $=-\frac{1}{2}\,{\operatorname{tr}\big(Q_{t}^{-1}\,e^{tB}\,Q\,e^{tB^{*}}\big)}+\frac{1}{2}\,\left|Q^{1/2}\,e^{tB^{*}}\,Q_{t}^{-1}\,(u-\,x)\right|^{2},$ (8.2)
$N_{t}^{(2)}(x,u)$ $=-\frac{1}{2}\,{\operatorname{tr}\big(Q_{t}^{-1}\,e^{tB}\,Q\,e^{tB^{*}}\big)}+\frac{1}{2t^{2}}\,\left|Q^{-1/2}(u-x)\right|^{2},$ (8.3)
$N_{t}^{(3)}(x,u)$ $=-\frac{n}{2t}+\frac{1}{2t^{2}}\,\left|Q^{-1/2}(u-x)\right|^{2}.$ (8.4)
Proof.
Differentiating (7.4), (7.5) and (7.6) with respect to $t$ and applying [13, Lemma 4.1], one easily obtains (8.1). ∎
In the following two lemmata we shall bound the $t$ derivative of $\widetilde{K}_{t}^{(\kappa-1)}(x,u)-\widetilde{K}_{t}^{(\kappa)}(x,u)$ for $\kappa=1,2,3$, and first write
$\partial_{t}\left(\widetilde{K}_{t}^{(\kappa-1)}(x,u)-\widetilde{K}_{t}^{(\kappa)}(x,u)\right)$ (8.5)
$=\,$ $\widetilde{K}_{t}^{(\kappa-1)}(x,u)\,\left(N_{t}^{(\kappa-1)}-N_{t}^{(\kappa)}\right)(x,u)+N_{t}^{(\kappa)}(x,u)\left(\widetilde{K}_{t}^{(\kappa-1)}-\widetilde{K}_{t}^{(\kappa)}\right)(x,u).$
We start by examining this difference for $\kappa=1$.
Lemma 8.2.
For $x,\>u\in 6B_{j}$ and $0<t\leq\min(1,|x_{j}|^{-2})$, one has
$\left|\partial_{t}\left(K_{t}(x,u)-\widetilde{K}_{t}^{(1)}(x,u)\right)\right|$ $\lesssim(1+|x|)\,\frac{e^{R(x)}}{t^{(n+1)/2}}\,\exp\left(-c\,\frac{|u-x|^{2}}{t}\right).$ (8.6)
Proof.
We first bound $N_{t}^{(0)}(x,u)-N_{t}^{(1)}(x,u)$, writing
$\Big|N_{t}^{(0)}(x,u)$ $-N_{t}^{(1)}(x,u)\Big|=\Big|\frac{1}{2}\,\left|Q^{1/2}\,e^{tB^{*}}\,Q_{t}^{-1}\,(u-D_{t}\,x)\right|^{2}$
$-\left\langle Q_{\infty}\,B^{*}\,Q_{\infty}^{-1}\,D_{t}\,x\,,\,(Q_{t}^{-1}-Q_{\infty}^{-1})\,(u-D_{t}\,x)\right\rangle-\frac{1}{2}\,\left|Q^{1/2}\,e^{tB^{*}}\,Q_{t}^{-1}\,(u-\,x)\right|^{2}\Big|$
$\lesssim\left|\left\langle Q_{\infty}\,B^{*}\,Q_{\infty}^{-1}\,D_{t}\,x\,,\,(Q_{t}^{-1}-Q_{\infty}^{-1})\,(u-D_{t}\,x)\right\rangle\right|$
$\qquad\quad+\Big|\frac{1}{2}\left|Q^{1/2}\,e^{tB^{*}}\,Q_{t}^{-1}\,(u-D_{t}\,x)\right|^{2}-\frac{1}{2}\left|Q^{1/2}\,e^{tB^{*}}\,Q_{t}^{-1}\,(u-\,x)\right|^{2}\Big|=:A+B.$
Using (3.4) and then (3.6), we obtain
$A$ $\lesssim\frac{|x|}{t}\,|u-D_{t}\,x|\lesssim\frac{|x|}{t}\left(|u-x|+|x-D_{t}\,x|\right)\lesssim\frac{|x|}{t}\,|u-x|+|x|^{2}.$
In a similar way,
$B$ $=\Big|\frac{1}{2}\left|Q^{1/2}\,e^{tB^{*}}\,Q_{t}^{-1}\,(u-x+x-D_{t}\,x)\right|^{2}-\frac{1}{2}\left|Q^{1/2}\,e^{tB^{*}}\,Q_{t}^{-1}\,(u-\,x)\right|^{2}\Big|$
$=\Big|\frac{1}{2}\left|Q^{1/2}\,e^{tB^{*}}\,Q_{t}^{-1}\,(x-D_{t}\,x)\right|^{2}+\langle Q^{1/2}\,e^{tB^{*}}\,Q_{t}^{-1}\,(u-x),Q^{1/2}\,e^{tB^{*}}\,Q_{t}^{-1}\,(x-D_{t}\,x)\rangle\Big|$
$\lesssim|x|^{2}+\frac{1}{t^{2}}\,|u-x|\,t|x|\lesssim|x|^{2}+\frac{|x|}{t}\,|u-x|,$
and thus
$\left|N_{t}^{(0)}(x,u)-N_{t}^{(1)}(x,u)\right|\lesssim|x|^{2}+\frac{|x|}{t}\,|u-x|.$
Hence, we have by (4.3)
$\left|K_{t}(x,u)\,\left(N_{t}^{(0)}(x,u)-N_{t}^{(1)}(x,u)\right)\right|$
$\lesssim\,\frac{e^{R(x)}}{t^{n/2}}\exp\left(-c\,\frac{|u-D_{t}\,x|^{2}}{t}\right)\left(|x|^{2}+\frac{|x|}{t}\,|u-x|\right)$
$\lesssim\,\frac{e^{R(x)}}{t^{n/2}}\,\exp\left(-c\,\frac{|u-x|^{2}}{t}\right)\,\exp(C\,{|u-x|\,|x|})\,\left(|x|^{2}+\frac{|x|}{t}\,|u-x|\right),$ (8.7)
the last step since
$|u-D_{t}\,x|^{2}$ $=|u-x|^{2}+2(u-x)\cdot(x-D_{t}\,x)+|x-D_{t}\,x|^{2}$
$\geq|u-x|^{2}-2C|u-x|\,t\,|x|.$
To estimate the last, polynomial factor in (8.1), we can replace |ux||u-x| by $\sqrt{t}$, if we also reduce the constant $c$ in the same formula. Further, $|x|^{2}\leq|x|/\sqrt{t}$ by our assumption on $t$, and so the polynomial factor is controlled by $|x|/\sqrt{t}$. The points $x$ and $u$ both have distance at most $6/(1+|x_{j}|)$ from $x_{j}$, which implies
$|u-x|\,|x|\lesssim(1+|x_{j}|)^{-1}\,(1+|x_{j}|)=1.$ (8.8)
Thus the last exponential factor in (8.1) is harmless, and we conclude that
$\left|K_{t}(x,u)\,\big(N_{t}^{(0)}(x,u)-N_{t}^{(1)}(x,u)\big)\right|\lesssim\frac{e^{R(x)}}{t^{n/2}}\,\exp\left(-c\,\frac{|u-x|^{2}}{t}\right)\,\frac{|x|}{\sqrt{t}}.$ (8.9)
Next we consider the term $N_{t}^{(1)}(x,u)\big(K_{t}(x,u)-\widetilde{K}_{t}^{(1)}(x,u)\big)$, and first observe that in view of (3.4)
$\big|N_{t}^{(1)}(x,u)\big|\lesssim\frac{1}{t}+\frac{|x-u|^{2}}{t^{2}}.$ (8.10)
We rewrite the long exponent in the expression (4.2) for $K_{t}(x,u)$ by replacing $u-D_{t}x$ by $u-x+x-D_{t}x$ and expanding the scalar product. This leads to
$-\frac{1}{2}\left\langle\left(Q_{t}^{-1}-Q_{\infty}^{-1}\right)(u-D_{t}x),u-D_{t}x\right\rangle=-\frac{1}{2}\left\langle\left(Q_{t}^{-1}-Q_{\infty}^{-1}\right)(u-x),u-x\right\rangle+E,$ (8.12)
where
$E$ $={-\left\langle\left(Q_{t}^{-1}-Q_{\infty}^{-1}\right)(u-x),x-D_{t}x\right\rangle}-\frac{1}{2}\left\langle\big(Q_{t}^{-1}-Q_{\infty}^{-1}\big)(x-D_{t}x),\,x-D_{t}x\right\rangle.$
From this and the expression (7.4) for $\widetilde{K}_{t}^{(1)}$, we see that we can factor out from $K_{t}(x,u)-\widetilde{K}_{t}^{(1)}(x,u)$ the exponential of the first right-hand term in (LABEL:KK), and get
$K_{t}(x,u)-\widetilde{K}_{t}^{(1)}(x,u)$ (8.14)
$=\left(\frac{\det\,Q_{\infty}}{\det\,Q_{t}}\right)^{{1}/{2}}e^{R(x)}\exp\Big(-\frac{1}{2}\left\langle(Q_{t}^{-1}-Q_{\infty}^{-1})(u-x),\,u-x\right\rangle\Big)\,(\exp E-1).$
We estimate $E$ by means of (3.6), (8.8) and the bound $t\lesssim 1/|x|^{2}$, obtaining
$|E|\lesssim\ |x|\,|u-x|+t|x|^{2}\lesssim 1.$
Thus
$|\exp E-1|\lesssim|E|\lesssim\ |x|\,|u-x|+t|x|^{2},$
and
$\left|K_{t}(x,u)-\widetilde{K}_{t}^{(1)}(x,u)\right|\lesssim\frac{e^{R(x)}}{t^{n/2}}\,\exp\left(-c\,\frac{|u-x|^{2}}{t}\right)\,\left(|x|\,|u-x|+t|x|^{2}\right).$ (8.15)
The polynomial factor here is $t$ times the one in (8.1), and arguing as there, we can replace it by $|x|\sqrt{t}$.
Combined with (8.10), this yields
$\Big|N_{t}^{(1)}(x,u)\Big(K_{t}(x,u)-\widetilde{K}_{t}^{(1)}(x,u)\Big)\Big|$ $\lesssim\frac{e^{R(x)}}{t^{n/2}}\,\exp\left(-c\,\frac{|u-x|^{2}}{t}\right)\,|x|\sqrt{t}\,\left(\frac{1}{t}+\frac{|x-u|^{2}}{t^{2}}\right)$
$\lesssim\frac{e^{R(x)}}{t^{n/2}}\,\exp\left(-c\,\frac{|u-x|^{2}}{t}\,\right)\,\frac{|x|}{\sqrt{t}}.$
Here we reduced the constant $c$ in order to replace |ux||u-x| by $\sqrt{t}$ in the polynomial factor.
This estimate and (8.9) imply Lemma 8.2. ∎
In the following lemma it is shown that the remaining two differences satisfy better bounds.
Lemma 8.3.
Let $x,\;u\in 6B_{j}$. Then for all $0<t\leq\min(1,|x_{j}|^{-2})$ one has
$\Big|\partial_{t}\Big(\widetilde{K}_{t}^{(\kappa-1)}(x,u)-\widetilde{K}_{t}^{(\kappa)}(x,u)\Big)\Big|$ $\lesssim\frac{e^{R(x)}}{t^{n/2}}\,\exp\left(-c\,\frac{|u-x|^{2}}{t}\right),\quad\kappa=2,3.\quad$
Proof.
We start with the case $\kappa=2$, using (8.5). By means of (3.2) and the fact that $e^{tB^{*}}=I_{n}(1+\mathcal{O}(t))$ as $t\to 0$, one finds that
$\left|N_{t}^{(1)}(x,u)-N_{t}^{(2)}(x,u)\right|$ $=\frac{1}{2}\,\Big|\big|Q^{1/2}\,e^{tB^{*}}\,Q_{t}^{-1}\,(u-\,x)\big|^{2}-\frac{1}{t^{2}}\,\left|Q^{-1/2}(u-x)\right|^{2}\Big|$
$\lesssim\frac{1}{t}\,|u-x|^{2}.$
Thus by (7.4) we have
$\left|\widetilde{K}_{t}^{(1)}(x,u)\,\big(N_{t}^{(1)}(x,u)-N_{t}^{(2)}(x,u)\big)\right|\lesssim\,\frac{e^{R(x)}}{t^{n/2}}\,\exp\left(-c\,\frac{|u-x|^{2}}{t}\right),$ (8.16)
for some $c>0$.
To consider the term $N_{t}^{(2)}(x,u)\big(\widetilde{K}_{t}^{(1)}(x,u)-\widetilde{K}_{t}^{(2)}(x,u)\big)$, we first notice that
$\left|N_{t}^{(2)}(x,u)\right|\lesssim\frac{1}{t}+\frac{|x-u|^{2}}{t^{2}}.$ (8.17)
Then we observe that (3.2) leads to
$\frac{1}{2}\langle(Q_{t}^{-1}-Q_{\infty}^{-1})(u-x),\,u-x\rangle=-\frac{1}{2t}\,|Q^{-1/2}(u-x)|^{2}+\mathcal{O}(|u-x|^{2}),\qquad t\to 0.$
This is analogous to (LABEL:KK), and arguing as in the steps leading from (LABEL:KK) to (8.15), we obtain since $|u-x|\lesssim 1$
$\left|\widetilde{K}_{t}^{(1)}(x,u)-\widetilde{K}_{t}^{(2)}(x,u)\right|\lesssim\frac{e^{R(x)}}{t^{n/2}}\,\exp\left(-c\,\frac{|u-x|^{2}}{t}\right)\,|u-x|^{2}.$
Combined with (8.17), this yields
$\Big|N_{t}^{(2)}(x,u)$ $\left(\widetilde{K}_{t}^{(1)}(x,u)-\widetilde{K}_{t}^{(2)}(x,u)\right)\Big|$
$\lesssim\frac{e^{R(x)}}{t^{n/2}}\,{\cancel{{e^{R(x)}}}}\exp\left(-c\,\frac{|u-x|^{2}}{t}\right)\,\left(\frac{|u-x|^{2}}{t}+\frac{|u-x|^{4}}{t^{2}}\right)\,$
$\lesssim\frac{e^{R(x)}}{t^{n/2}}\,\exp\left(-c\,\frac{|u-x|^{2}}{t}\right),$
with a modified value of $c$.
The estimate of the lemma for $\kappa=2$ follows from this and (8.16).
Moving to $\kappa=3$, we observe by (3.2) that
$\left|N_{t}^{(2)}(x,u)-N_{t}^{(3)}(x,u)\right|$ $=\Big|-\frac{1}{2}\,{\operatorname{tr}\big(Q_{t}^{-1}\,e^{tB}\,Q\,e^{tB^{*}}\big)}+\frac{n}{2t}\Big|$
$=\left|-\frac{1}{2t}\,{\operatorname{tr}\big(\,I_{n}+\mathcal{O}(t)\big)}+\frac{n}{2t}\right|=\mathcal{O}(1),$
and (7.5) implies
$\left|\widetilde{K}_{t}^{(2)}(x,u)\left(N_{t}^{(2)}(x,u)-N_{t}^{(3)}(x,u)\right)\right|\lesssim\frac{e^{R(x)}}{t^{n/2}}\exp\left(-c\,\frac{|u-x|^{2}}{t}\right).$ (8.18)
Then we consider the term $N_{t}^{(3)}(x,u)\left(\widetilde{K}_{t}^{(2)}(x,u)-\widetilde{K}_{t}^{(3)}(x,u)\right)$ and observe that
$\left|N_{t}^{(3)}(x,u)\right|\lesssim\frac{1}{t}+\frac{|x-u|^{2}}{t^{2}}.$ (8.19)
From (3.3) we see that
$\left|\widetilde{K}_{t}^{(2)}(x,u)-\widetilde{K}_{t}^{(3)}(x,u)\right|$
$=\big({\det Q_{\infty}}\big)^{1/2}e^{R(x)}\exp\left(-\frac{1}{2t}\,\left|Q^{-1/2}(u-x)\right|^{2}\right)\left|\frac{1}{{(\det Q_{t})}^{1/2}}-\frac{1}{(\det Q)^{1/2}\,t^{n/2}}\right|$
$\lesssim\frac{e^{R(x)}}{t^{n/2}}\,\exp\Big(-\frac{1}{2t}\,|u-x|^{2}\Big)\,t.$
With (8.19) we then have
$\Big|N_{t}^{(3)}(x,u)\,$ $\Big(\widetilde{K}_{t}^{(2)}(x,u)-\widetilde{K}_{t}^{(3)}(x,u)\Big)\Big|$
$\lesssim\frac{e^{R(x)}}{\,t^{n/2}}\,\exp\left(-\frac{1}{2t}\,|u-x|^{2}\right)\,\left(1+\frac{|x-u|^{2}}{t}\right)\,$
$\lesssim\frac{e^{R(x)}}{t^{n/2}}\,\exp\left(-c\,\frac{|u-x|^{2}}{t}\right),$
Together with (8.18), this proves the lemma for $\kappa=3$. ∎
8.2. Boundedness of the variation operator for the differences
By means of the estimates proved in the previous subsection, we can finally estimate the variation of $\Delta_{t}^{(\kappa)}\,f$, for $\kappa=1,2,3$.
Proposition 8.4.
Let $1\leq\varrho<\infty$ and $j\in\mathbb{N}$. The operator that maps $f\in L^{1}(\gamma_{\infty})$ to the function
$x\mapsto\|\widetilde{r}_{j}(x)\,\Delta_{t}^{(\kappa)}\,(f\,r_{j})(x)\|_{v(\varrho),(0,\min(1,|x_{j}|^{-2})]},\qquad\;\kappa=1,2,3,$
is of strong type $(p,p)$ for all $p\geq 1$ with respect to the invariant (and Lebesgue) measure. This is uniform in $j$.
Because of (1.6) and (1.5), this proposition implies
$J_{2}^{1,\infty}\left(\widetilde{r}_{j}\,\Delta_{t}^{(\kappa)}\,(f\,r_{j}):\,0<t\leq\min(1,|x_{j}|^{-2})\right)\lesssim\|f\|_{L^{1}(4B_{j},\gamma_{\infty})},\qquad\;\kappa=1,2,3,$ (8.20)
and for $1<p<\infty$
$J_{2}^{p}\left(\widetilde{r}_{j}\,\Delta_{t}^{(\kappa)}\,(f\,r_{j}):\,0<t\leq\min(1,|x_{j}|^{-2})\right)\lesssim\|f\|_{L^{p}(4B_{j},\gamma_{\infty})},\qquad\;\kappa=1,2,3.$ (8.21)
Proof of Proposition 8.4.
We start with the case of $\Delta_{t}^{(1)}$. Let $f\in L^{1}(\gamma_{\infty})$. We want to apply an argument like (4.6) to the operator $f\mapsto\widetilde{r}_{j}\,\Delta_{t}^{(1)}\,(f\,r_{j})$ and its kernel
$\widetilde{r}_{j}(x)\,\left(K_{t}(x,u)-\widetilde{K}_{t}^{(1)}(x,u)\right)\,r_{j}(u).$
What must then be verified is only the swap of differentiation and integration in the last step of (4.6), that is,
$\frac{\partial}{\partial t}\,\int\widetilde{r}_{j}(x)\,\left(K_{t}(x,u)-\widetilde{K}_{t}^{(1)}(x,u)\right)\,r_{j}(u)\,f(u)\,d\gamma_{\infty}(u)$
$=$ $\>\int\widetilde{r}_{j}(x)\,\frac{\partial}{\partial t}\,\left(K_{t}(x,u)-\widetilde{K}_{t}^{(1)}(x,u)\right)\,r_{j}(u)\,f(u)\,d\gamma_{\infty}(u)$ (8.22)
for $0<t<\min(1,|x_{j}|^{-2}).$ For this, we choose two points $t_{0}<t$ in the interval $\left(0,\,\min(1,|x_{j}|^{-2})\right)$ and replace $t$ by $\tau$ in the right-hand side of (8.2), which we then integrate $d\tau$ over $t_{0}<\tau<t.$ We get
$\int_{t_{0}}^{t}\,\int\,\widetilde{r}_{j}(x)\,\frac{\partial}{\partial\tau}\,\left(K_{\tau}(x,u)-\widetilde{K}_{\tau}^{(1)}(x,u)\right)\,r_{j}(u)\,f(u)\,d\gamma_{\infty}(u)\,d\tau$ (8.23)
$=$ $\>\int\,\int_{t_{0}}^{t}\,\widetilde{r}_{j}(x)\,\frac{\partial}{\partial\tau}\,\left(K_{\tau}(x,u)-\widetilde{K}_{\tau}^{(1)}(x,u)\right)\,r_{j}(u)\,f(u)\,d\tau\,d\gamma_{\infty}(u),$ (8.24)
the equality by means of Fubini’s theorem. To see that Fubini can be applied here, we observe that Lemma 8.2 implies
$\int_{t_{0}}^{t}\,\int\widetilde{r}_{j}(x)\,\left|\frac{\partial}{\partial\tau}\,\left(K_{\tau}(x,u)-\widetilde{K}_{\tau}^{(1)}(x,u)\right)\right|\,r_{j}(u)\,|f(u)|\,d\gamma_{\infty}(u)\,d\tau$
$\lesssim$ $\>(1+|x|)\,e^{R(x)}\,\int_{t_{0}}^{t}\frac{d\tau}{\tau^{(n+1)/2}}\,\int|f(u)|\,d\gamma_{\infty}(u)<\infty.$
Having thus verified (8.23), we can rewrite its right-hand side by a trivial evaluation of the inner integral, and obtain
$\int_{t_{0}}^{t}\,\int\widetilde{r}_{j}(x)\,\frac{\partial}{\partial\tau}\,\left(K_{\tau}(x,u)-\widetilde{K}_{\tau}^{(1)}(x,u)\right)\,r_{j}(u)\,f(u)\,d\gamma_{\infty}(u)\,d\tau$ (8.25)
$=$ $\>\int\widetilde{r}_{j}(x)\,\big(K_{t}(x,u)-\widetilde{K}_{t}^{(1)}(x,u)\big)\,r_{j}(u)\,f(u)\,d\gamma_{\infty}(u)$ (8.26)
- $\>\int\widetilde{r}_{j}(x)\,\big(K_{t_{0}}(x,u)-\widetilde{K}_{t_{0}}^{(1)}(x,u)\big)\,r_{j}(u)\,f(u)\,d\gamma_{\infty}(u).$ (8.27)
Differentiating this equation with respect to $t$, we obtain (8.2) as desired.
We can thus apply (4.6) with the indicated kernel. Together with Lemma 8.2, this yields
$\|\widetilde{r}_{j}(x)\,\Delta_{t}^{(1)}(f\,r_{j})(x)\|_{v(\varrho),(0,\min(1,|x_{j}|^{-2})]}$ (8.28)
$\lesssim$ $\int\,\int_{0}^{\min(1,|x_{j}|^{-2})}\,\widetilde{r}_{j}(x)\,\left|\frac{\partial}{\partial t}\,\left(K_{\tau}(x,u)-\widetilde{K}_{\tau}^{(1)}(x,u)\right)\right|\,r_{j}(u)\,dt\,|f(u)|\,d\gamma_{\infty}(u)$ (8.29)
$\lesssim$ $\,(1+|x|)\,\,e^{R(x)}\int\int_{0}^{\min(1,|x_{j}|^{-2})}t^{-(n+1)/2}\,e^{-c\,\frac{|u-x|^{2}}{t}}\,dt\,\,\,|f(u)|\,r_{j}(u)\,d\gamma_{\infty}(u)\,\mathbbm{1}_{6B_{j}}(x).$ (8.30)
In the inner integral here, we make the transformation $s=|u-x|^{2}/t$ and find
$\int_{0}^{\min(1,|x_{j}|^{-2})}t^{-(n+1)/2}\,\exp\big(-c\,\frac{|u-x|^{2}}{t}\big)\,dt$ (8.31)
$=$ $\int_{|u-x|^{2}/\min(1,|x_{j}|^{-2})}^{\infty}|u-x|^{1-n}\,s^{\frac{n-3}{2}}\,e^{-cs}\,ds.$ (8.32)
If $n>1$, the last integral is at most $C\,|u-x|^{1-n}$. But if $n=1$, it is bounded by
$C+\log_{+}\frac{\min(1,|x_{j}|^{-2})}{|u-x|^{2}},$
where $\log_{+}\beta=\max(0,\beta)$ for $\beta>0$. We insert these estimates in the last double integral in (8.28), after observing that there $u-x\in{4B_{j}}-{6B_{j}}\subset B(0,10/(1+|x_{j}|))$ and also $e^{R(u)}\simeq e^{R(x)}\simeq e^{R(x_{j})}$. Thus we can replace $d\gamma_{\infty}(u)$ by $du$ and delete the factor $e^{R(x)}$, in this integral. Further, $1+|x|\lesssim 1+|x_{j}|$ there. For $n>1$ we get
$\|\widetilde{r}_{j}(x)\,\Delta_{t}^{(1)}(f\,r_{j})(x)\|_{v(\varrho),(0,\min(1,|x_{j}|^{-2})]}\lesssim(1+|x_{j}|)\,\int\,|u-x|^{1-n}\,|f(u)|\,r_{j}(u)\,du\,\mathbbm{1}_{6B_{j}}(x).$
The right-hand side here is essentially a convolution, and we can estimate it by
$|f\,r_{j}|*g(x),$
where $g(y)=C\,(1+|x_{j}|)\,|y|^{1-n}\,\mathbbm{1}_{B(0,10/(1+|x_{j}|))}(y)$, when $n>1.$
For $n=1$ we arrive at a similar convolution, but now
$g(y)=(1+|x_{j}|)\,\left(C+2\log_{+}\frac{\min(1,|x_{j}|^{-1})}{|y|}\right)\,\mathbbm{1}_{B(0,10/(1+|x_{j}|))}(y).$
In both cases, we find that $\|g\|_{L^{1}(dy)}\lesssim 1$, and so convolution by $g$ defines a bounded operator on all $L^{p}(dx)$ spaces.
This proves Proposition 8.4 for $\kappa=1$.
The cases $\kappa=2,\;3$ can be done in exactly the same way, since, as already observed, Lemma 8.3 gives a stronger estimate than Lemma 8.2.
The proposition is proved. ∎
9. The local case for small $t$ : the main operator $M_{t}$
We are left with the estimate of the main part $M_{t}f$ in (7.1), that is, the integral operator with the kernel $\widetilde{K}_{t}^{(3)}$ defined in (7.6).
Proposition 9.1.
Let $j\in\mathbb{N}$. For all $p>1$ one has
$J_{2}^{p}\left(\widetilde{r}_{j}\,M_{t}(f\,r_{j}):0<t\leq\min(1,|x_{j}|^{-2})\right)\lesssim_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p}\|f\,r_{j}\|_{L^{p}(\gamma_{\infty})}$ (9.1)
and, moreover,
$J_{2}^{1,\infty}\left(\widetilde{r}_{j}\,M_{t}(f\,r_{j}):0<t\leq\min(1,|x_{j}|^{-2})\right)\lesssim\|f\,r_{j}\|_{L^{1}(\gamma_{\infty})}.$ (9.2)
The implicit constants in these inequalities are uniform in $j$.
Proof.
Let for $t>0$
$\psi_{t}(y)=\Big(\frac{\det Q_{\infty}}{\det Q}\Big)^{1/2}\,{t^{-n/2}}\,\exp\big(-\frac{1}{2t}\,|Q^{-1/2}\,y|^{2}\big),\qquad y\in\mathbb{R}^{n}.$
Then
$M_{t}(f\,r_{j})(x)$ $=e^{R(x)}\Big(\psi_{t}*\big(fr_{j}e^{-R(\cdot)}\big)\Big)(x).$
for $f\in L^{1}(\gamma_{\infty})$. We recall that $e^{R(x)}\simeq e^{R(x_{j})}$ for $x\in 6B_{j}\supset\mathrm{supp}\,\widetilde{r}_{j}$.
If $\widetilde{r}_{j}(x)\,M_{t}(f\,r_{j})(x)=\widetilde{r}_{j}(x)\,e^{R(x)}\left(\psi_{t}*\big(fr_{j}e^{-R(\cdot)}\big)\right)(x)$ has a jump of amplitude larger than $\lambda$ for some $x\in\mathrm{supp}\,\widetilde{r}_{j}$, then $\psi_{t}*\big(fr_{j}e^{-R(\cdot)}\big)(x)$ has a jump of amplitude larger than $c\,\lambda\,e^{-R(x_{j})}$ for some $c>0$. As before, the jumps are in the variable $t$. Thus
${N_{\lambda}\big(\widetilde{r}_{j}(x)\,M_{t}(f\,r_{j})(x)}\big)\leq N_{c\,\lambda e^{-R(x_{j})}}\left(\psi_{t}*\big(fr_{j}e^{-R(\cdot)}\big)\right)(x)\,\mathbbm{1}_{6B_{j}}(x),$ (9.3)
and for all $\alpha>0$,
$\gamma_{\infty}\left\{x\in 6B_{j}:\,\lambda\,\sqrt{N_{\lambda}\left(\widetilde{r}_{j}(x)\,M_{t}(f\,r_{j})\right)(x)}>\alpha\right\}$
$\lesssim\gamma_{\infty}\left\{x\in 6B_{j}:\,\lambda\,\sqrt{N_{c\lambda e^{-R(x_{j})}}\,\left(\psi_{t}*\big(fr_{j}e^{-R(\cdot)}\big)\right)(x)}>\alpha\right\}$
$=\gamma_{\infty}\left\{x\in 6B_{j}:\,c\,e^{-R(x_{j})}\,\lambda\,\sqrt{N_{c\lambda e^{-R(x_{j})}}\,\left(\psi_{t}*\big(fr_{j}e^{-R(\cdot)}\big)\right)(x)}>c\,e^{-R(x_{j})}\,\alpha\,\right\}$
$\simeq e^{-R(x_{j})}\left|\left\{x\in 6B_{j}:\,ce^{-R(x_{j})}\,\lambda\,\sqrt{N_{c\lambda e^{-R(x_{j})}}\,\left(\psi_{t}*\big(fr_{j}e^{-R(\cdot)}\big)\right)(x)}>\,c\,e^{-R(x_{j})}\,\alpha\right\}\right|.$
Now we apply Theorem 2.4 in [28] to the last expression here, getting
$\gamma_{\infty}$ $\Big\{x\in 6B_{j}\,:\,\lambda\,\sqrt{N_{\lambda}\left(\widetilde{r}_{j}(x)\,M_{t}(f\,r_{j})\right)(x)}>\alpha\Big\}$
$\lesssim$ $\,\frac{e^{-R(x_{j})}}{e^{-R(x_{j})}\,\alpha}\,\left\|\psi_{t}*\big(fr_{j}e^{-R(\cdot)}\big)\right\|_{L^{1}(du)}\lesssim\,\frac{1}{\alpha}\,\|f\,r_{j}\,e^{-R(\cdot)}\|_{L^{1}(du)}=\frac{1}{\alpha}\,\|f\,r_{j}\|_{L^{1}(\gamma_{\infty})},$
where we also used the fact that convolution with $\psi_{t}$ defines a bounded operator on $L^{1}(du)$. This is uniform in $\lambda>0$ and in $j$, and proves (9.2).
We now consider the strong type $(p,p)$ for $1<p<\infty$. Starting from (9.3), moving to Lebesgue measure and then applying [28, Theorem 2.4], one finds
$\left\lVert\,\lambda\sqrt{N_{\lambda}\big(\widetilde{r}_{j}(\cdot)\,M_{t}(f\,r_{j})\big)(\cdot)}\,\right\rVert_{L^{p}(6B_{j},\gamma_{\infty})}\leq\left\lVert\,\lambda\sqrt{N_{c\,\lambda e^{-R(x_{j})}}\left(\big(\psi_{t}*\big(fr_{j}e^{-R(\cdot)}\big)\right)}\right\rVert_{L^{p}(6B_{j},\gamma_{\infty})}$
$=c^{-1}e^{R(x_{j})}\left\lVert\,ce^{-R(x_{j})}\lambda\sqrt{N_{c\,\lambda e^{-R(x_{j})}}\left(\big(\psi_{t}*\big(fr_{j}e^{-R(\cdot)}\big)\right)}\,\right\rVert_{L^{p}(6B_{j},\gamma_{\infty})}$
$\simeq_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p}e^{R(x_{j})}\,e^{-R(x_{j})/p}\left\lVert\,ce^{-R(x_{j})}\lambda\sqrt{N_{c\,\lambda e^{-R(x_{j})}}\left(\big(\psi_{t}*\big(fr_{j}e^{-R(\cdot)}\big)\right)}\,\right\rVert_{L^{p}(6B_{j},dx)}$
$\lesssim_{p}e^{R(x_{j})}\,e^{-R(x_{j})/p}\left(\int\big|f(u)\,r_{j}(u)\,e^{-R(u)}\big|^{p}du\right)^{1/p}\simeq_{p}\|f\,r_{j}\|_{L^{p}(\gamma_{\infty})},$
uniformly in $\lambda>0$ and $j$. This is (9.1), and Proposition 9.1 is proved. ∎
Now the remaining part of Theorem 5.2 with $t\in(0,\min(1,|x_{j}|^{-2})]$ follows as a combination of (8.20) and (9.2). Theorem 1.2 is completely proved.